Find the given derivative by finding the first few derivatives and observing the pattern that occurs. 298 (sin(x)) 0x 98
Answer #1

98 Given d ( sinx) : ? 98 da scx) = cos(x) of CX) = -siny fCX 2 = COS X of (x) = sin(x) 4 will result in sin(x). so; e
Answer #2

SOLUTION :


d/dx (sin (x)) = cos(x)


=> d2/dx2  (sin (x)) = d/dx (cos (x)) = - sin (x)


=> d3/dx3  (sin (x)) = d/dx (- sin (x)) = - cos  (x)


=> d4/dx4  (- sin (x)) = d/dx (- cos (x)) = sin (x)


Every 4th derivative = sin(x)


So,


d96/dx96 (sin (x)) = sin (x)


=> d97/dx97 (sin (x)) = cos (x)


=> d98/dx98 (sin (x)) = - sin (x)  (ANSWER).

answered by: Tulsiram Garg